OpenStudy (anonymous):

WORD PHRASE TO ALGEBRAIC EXPRESION FOUR MORE THAN TWICE A NUMBER IS 5

OpenStudy (anonymous):

Okay so 4 = Four More = + Twice = *2 A number is an exponent So im guessing its 4 + a*2 = 5?

OpenStudy (anonymous):

I said exponent I meant variable sorry.

OpenStudy (anonymous):

Do you get what I am saying?

OpenStudy (texaschic101):

2a + 4 = 5

OpenStudy (anonymous):

same thing but diffrent order you know lol

OpenStudy (anonymous):

4+2a=5, or 2a-1